OpenStudy (anonymous):

Calculate the density of a rectangular piece of wood from the following data: Length: 25 cm. Width: 10 cm. Height: 4 cm. Mass: 15 g

OpenStudy (souvik):

volume v=25*10*4=1000 cm^3 mass m=15g density D=M/D=15/1000=0.015 g/cm^3